Nature’s Treasure tree farm is trying to decide which fertilizer to use to grow its trees. It can either buy the famous Greenwood Fertilizer, which costs $5 and can grow trees which can be sold for $8 in 4 years, or it can buy Brian and Peter’s Fertilizer which costs $4 and can grow trees which can be sold for $8 in 6 years. What is the rate of return of each option?
Answer #1

Rate of return of each option:-

Greenwood fertilizer

=(1+r)^4 = 8/5

=(1+r)^4 = 1.6

= (1+r) = 1.6^1/4

= (1+r) = 1.124683

so r = 1.124683-1 = 0.1246 = 12.46%

Brian and Peter's

= (1+r)^6 = 8/4

= (1+r)^6 = 2

= (1+r) = 2^1/6

= (1+r) = 1.1224

r = 1.1224-1 = 0.1224 = 12.24%